A satellite has a mass of 6392 kg and is in a circular orbit 4.88 × 105 m above the surface of a planet. The period of the orbit is 1.7 hours. The radius of the planet is 4.89 × 106 m. What would be the true weight of the satellite if it were at rest on the planet’s surface?
Answer #1

First of all, we need to find the mass of the planet. For that we will use Kepler's law

M = 42(r3) / (T2G)

M = 42 (4.89e6 + 4.88e5)3 / ( 1.72 * 36002 * 6.67e-11)

M = 2.458e24 Kg

Now, we can find value of g for the planet

g = GM / r2

g = 6.67e-11 * 2.458e24 / 4.89e62

g = 6.856 m/s2

therefore,

true weight on the surface

W = mg

W = 6392 * 6.856

W = 43826.65 N